Calculate the value of dE and dH on heating 64.0 grams of oxygen from 273 k to 373 k . Cv and Cp on an average are 5 and 7 cal/mol/deg respectively

Grade:12th pass

1 Answers

Shweta Ahire
5 years ago
Delta E = heat(q) exchanged at constant volume 
So q= nCv deltaT
        = 64/32 × 5 ×100
    Delta  E =1000 cal =4.3 kJ
 
 delta H = heat exchanged at constant pressure 
So q = nCp deltaT
         =64/32 × 7 × 100
      Delta H  = 1400 cal = 5.88kJ